OpenStudy (anonymous):

NEED HELP!! WILL FAN AND GIVE YOU ALOT OF MEDALS!!!!!! 1.simplify the radical expression √64x^8 A.8x B.8x^4 C.8√x^8 D.8x^2 2.simplify the rational expression. √125h^4 A.5h^2√5 B.25h^2√5 C.h√62.5 D.5√5h^4 3.multiply. give your answer in simplest form. 2√9*3√15 A.18√15 B.6√6 C.6√24 D.18√6

OpenStudy (anonymous):

1. How would you convert a square root into a fractional index?

OpenStudy (anonymous):

@summertime

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

none at all! im so confused!!

OpenStudy (anonymous):

@AnImEfReaK

OpenStudy (anonymous):

Okay. You need to know this index law. \(\sqrt{a}\) \(=\) \(a^\dfrac{1}{2}\) When a number is under a square root, you can turn the square root into a power of \(\dfrac{1}{2}\).

OpenStudy (anonymous):

This applies to a long number with many terms in it.